# Assessing realistic binding energies of some essential interstellar radicals with amorphous solid water:

### v1

@@ 3.3 Comparison of the BEs with previous studies
Open-shell species pose significant computational challenges due to multireference characteristics and the need to account for static and dynamic correlation effects. Accurately capturing the true energy of these species requires a sensitive and careful approach, often involving multireference wave function theories (e.g., coupled cluster method) and perturbation methods (e.g., Møller-Plesset method). Although these techniques can provide highly accurate results, they come with a substantial computational cost. The combination of the B97X-D functional and the 6-311+G(d,p) basis set offers a strong solution to achieve a practical balance between accuracy and efficiency. The B97X-D functional is crucial for handling the complex electronic structures inherent to radicals and is a versatile tool for studying a wide range of systems. The 6-311+G(d,p) basis set, with its valence triple- quality and added polarization functions, further enhances the accuracy by providing a flexible and detailed description of electron density, especially where electron distribution is uneven due to unpaired electrons. This combination is not necessarily the best or most accurate option available. Still, it is a well-considered choice that delivers a high degree of accuracy while keeping computational costs relatively low. Furthermore, we believe our methodology with careful structural optimization and analysis is robust enough to be applied to other radical species and, more generally, to species with challenging electronic structures.
@@ 4.1 Chemisorption
Some of our binding sites exhibit chemisorption for CH (a chemical bond is formed with one oxygen atom of a water molecule of the cluster). Figure 9 depicts the formation of CH2OH and CHOH2 by chemisorption. There is a possibility that the chemisorbed CH2OH could produce CH3OH by the Eley-Rideal mechanism by capturing an H atom from the gas phase. Alternatively, a chemisorbed H atom could react with CH2OH to form methanol. Recent experiments by Grieco et al. (2024) have suggested the formation of H2O at a high temperature (approximately 85 K) by chemisorbed H and O. Additionally, Grieco et al. (2023) reported the formation of H2 up to a temperature as high as 250 K, suggesting the feasibility of the hydrogenation reaction at higher temperatures than previously thought. Contrary to previous understanding of methanol formation through the surface hydrogenation of CO, Molpeceres et al. (2021) and Tsuge et al. (2023) suggest that before CO freezeout, atomic carbon deposited on ASW could form H2CO. Tsuge et al. (2023) pointed out that at temperatures below 20 K, some of the accreted carbon atoms are quickly converted to formaldehyde. The remaining physisorbed atoms are partially changed to chemisorbed and hydrogenated to CH4. Physisorbed carbon atoms can diffuse at temperatures between 20 K and 30 K, promoting insertion reactions with other carbon atoms or species. Future observations with the James Webb Space Telescope (JWST) could provide additional details about the significance of chemisorption in forming COMs in the ice phase.
4.2 Implication of the computed BEs on Astrochemical model
@@
Recently, using ALMA observations, Coutens et al. (2019) detected multiple transitions of HONO in the low-mass protostellar system IRAS 16293–2422. This system consists of three protostars, two located toward source A and one toward source B (Maureira et al., 2020). The identification by Coutens et al. (2019) specifically focuses on the protostar associated with source B. They also attempted to explain the abundances of HONO-related species and their ratios using a chemical model. However, they were unable to account for all the observed features. In their study, Coutens et al. (2019) estimated abundance ratios of HONO/NH2OH and HONO/HNO to be greater than or equal to and , respectively. Using their chemical model, they successfully explained the abundance of HONO, N2O, and NO2. However, their model resulted in an overproduction of NH2OH and HNO, leading to lower HONO/NH2OH () and HONO/HNO (0.01) ratios compared to the observed values. Our study suggests a weak binding of NO with the water substrate. It would be interesting to see how the estimated BEs would affect the modeled abundances of these species.
@@
We utilize the Rokko code Furuya et al. (2015, 2017) for our chemical modeling, which encompasses three phases: gas, icy grain surface, and homogeneous bulk ice mantle. Our active ice layers are limited to the top four monolayers. All layers, including the active ice layers, are considered part of the bulk ice mantle. The gas and ice phase networks are based on Garrod (2013). The active layers undergo photodissociation and photodesorption of species. A constant reactive desorption of all the species is assumed with a factor of 0.01, resulting in a single product. In molecular dynamics studies, it had been determined that approximately 10% of the formation energy is retained after the formation of H2S from HS+H (Bariosco et al., 2024, and references therein). However, based on experimentally determined reaction desorption probabilities for H2S, HDS, and D2S (Oba et al., 2019; Furuya et al., 2022), a reactive desorption factor of 0.03 is being utilized. A standard cosmic-ray ionization rate of s-1 is considered.
@@ 4.2.1 Dark Cloud
For the dark cloud model, we start with a hydrogen number density, cm-3, visual extinction () of 10 mag, and a gas and ice temperature of 10 K. The initial elemental abundances of species are taken from Garrod et al. (2017). The upper panel of Fig. 5 shows the time evolution of the radicals. Here, “G” refers to the notation when the whole BE sets are from Garrod (2013), whereas “G+S” refers to the BEs for the radicals taken from this work, and the rest are considered from Garrod (2013). In this work, we are using a default diffusion-to-BE ratio of 0.4. Since no estimated BEs were available for HONO, H2NO2, and H3NO2, we used our estimated values of 2918 K, 4205 K, and 3044 K, respectively, in both “G” and “G+S.” It shows that the changes in the BE would have a minor effect on the gas phase abundances of these radicals. However, it shows a major effect on the ice phase abundances of these radicals, especially for the ice phase NO and CH. NO ice abundance is almost reduced by a factor of at years. Meanwhile, the ice phase abundance of CH increased by a factor of 100. The reduction in the abundance of NO is due to the weak binding of NO obtained with this work ( K) compared to the earlier estimate (1600 K; Wakelam et al., 2017). The increase in the surface abundance of CH radicals is due to the estimate of a higher BE ( K) than previous work (Garrod, 2013, 925 K). The shaded lines on the graph indicate the uncertainty of our BE values from Table 2. Since NO has a very weak binding with the water substrate, its small uncertainty (about 94 K) in the BE estimation could cause a significant change in its ice phase abundance (see the black shaded area in the top right panel of Fig. 5).
The lower panel of Fig. 5 shows the time evolution of the NO-related species. It clearly shows that with the present BE estimation of NO, we could expect a favorable formation of gas phase HONO in a dark cloud. On the contrary, the formation of gas phase NH2OH is heavily affected, which could be useful in explaining the observed HONO/NH2OH ratio in star-forming regions (Coutens et al., 2019) discussed in 4.2.2. The shaded regions represent when the uncertainty in the BE calculations is considered. The obtained abundances and abundance ratio with HONO are noted in Table 4. Due to the weak binding of NO obtained in this study, we have a higher abundance of the ice phase NO2 through the reaction: , resulting in a favorable formation of HONO-gr by . The formation of NH2OH-gr process through successive hydrogenation reactions: . When our BE value of NO is used, due to the high abundance of atomic oxygen on the grain compared to the atomic hydrogen, NO-gr is mainly channelized to form , resulting in a decrease in the formation of . On the contrary, it facilitates the HONO-gr formation. In the lower panel of Fig. 5, we have observed significant uncertainties in the ice phase abundances of HNO and NH2OH due to the uncertainty surrounding the BE of NO ( 94 K).
@@
s when the BEs calculated here are used and for others Garrod (2013) used.
@@ 4.2.2 Hot corino
We have further examined a low-mass star-forming region to assess the impact of our proposed BEs on abundance estimation. This study considers three stages: the isothermal collapsing stage, the warm-up stage, and the post-warm-up stage. During the isothermal collapsing stage, the cloud is allowed to collapse from a minimum density of cm-3 to a maximum density of cm-3 in one million years, while maintaining a gas temperature of 10 K and ice temperature of 10 K. The initial abundance for this model is taken from the final abundance of the dark cloud model discussed in 4.2.1. The visual extinction at the first stage can vary with the density following the relation provided by Garrod & Pauly (2011), resulting in a variation of . In the warm-up stage, the density and visual extinction parameters are consistently set to their maximum values for an additional one million years while the temperature gradually rises to a peak of K. During the post-warm-up stage, the parameters listed above remain at their maximum values for an additional years.
@@
The upper panel of Fig. 6 shows the time evolution of the abundances of gas-phase radicals. With the new BE value, the gas-phase abundance of NO (black dashed curve) has decreased significantly during the warm-up stage. In the lower panel of Fig. 6, we show the abundances of NO and its related species separately. The uncertainties in the BE calculations are also considered here as it was discussed in the context of Fig. 5. In Table 4, we compare the peak abundances (peak value is taken beyond the isothermal stage) and the peak abundance ratio of some NO-related species obtained from our simulations with that observed by Coutens et al. (2019). With the new BEs (represented by the dashed curves), we observe a significant reduction in the abundances of NO, N2O, HNO, and NH2OH. Conversely, there is a surge in the abundance of HONO.
@@
NH2OH is primarily produced in the cold ice phase through the successive sequential hydrogenation of NO-gr, as discussed in 4.2.1, and it can sublimate to the gas phase after around years, depending on its BE considered (6806 K; Garrod, 2013). Due to the weak binding of NO, its abundance in the ice phase decreases as it forms NO2, leading to a decrease in NH2OH. We estimated a BE of 2918 K for HONO, so it is expected to be released earlier than NH2OH in the gas phase. However, an additional radical-radical pathway () is responsible for ice phase HONO production in warmer regions. The low BE of NO obtained here further facilitates this radical-radical ice phase reaction.
@@
In our study, with our estimated BE of NO, we achieve a lower abundance of NH2OH and HNO (see Fig. 6 and Table 4). We obtain peak abundance ratios of HONO/NH2OH and HONO/HNO of 143 and 3781, respectively, consistent with the observation. Table 4 depicts that the other ratios from the observations seem to be between the previous model (Coutens et al., 2019) and this model. This may indicate that the diffusion activation energy of NO was overestimated in the previous one and underestimated in the present.
@@ 5 Conclusions
This study aims to calculate the BE of certain interstellar radicals on cluster surfaces of ASW. The work employs a fully quantum chemical approach to examine a large number of binding sites for all the considered radicals using a suitable ASW cluster model. The significant findings of our work are presented below:
@@
We estimate the binding energies (BEs) of 7 diatomic radicals using a pure quantum chemical approach with high-level water clusters and methods, and discuss the implications of the computed BEs on astrochemical models.
@@
In our study, the BSSE correction is carefully employed extensively using the CP method. This correction is crucial, especially with small and medium-sized basis sets, for obtaining accurate and reliable estimates of the BEs between radical species and the ASW cluster.
@@
Our results suggest that most of the time, we have physisorption, indicating weak bonding between radicals and the water molecule. However, in addition to physisorption, we find chemisorption of CH with cluster in two cases. In this case, a chemical bond is formed between CH and an oxygen atom of a water molecule. Figure 9 illustrates the production of CH2OH through the chemisorption of CH on a water substrate. This warrants the potential formation of methanol through hydrogenation in warmer regions. Future observations with the JWST could address this issue. Moreover, our estimated value of the BE (physisorption) of CH is 2-3 times higher than the available guess value. Hence, we suggest using these values for astrochemical models.
@@
We obtain a significantly lower BE value of NO (704 K), which is in excellent agreement with a very recent study (average value of 712 K with ASW model; Martínez-Bachs et al., 2024) compared to that used earlier (1600 K). Our obtained BEs of NO would be useful in explaining the observed HONO/NH2OH and HONO/HNO ratio obtained in the B component of the low-mass protostellar binary IRAS 16293–2422.
@@ Acknowledgements.
The authors thank the referee for the valuable insights and constructive comments that significantly improved the manuscript. This work is completed while M.S. is a postdoctoral fellow at the Institute of Planetology and Astrophysics of Grenoble (IPAG), University of Grenoble Alpes. M.S. acknowledges financial support through the European Research Council (consolidated grant COLLEXISM, grant agreement ID: 811363). N.I. acknowledges the support of FONDECYT Regular Grant 1241193 and Vicerectoría de Investigación y Postgrado (VRIP). A.D. acknowledges the support of MPE for sponsoring a scientific visit to MPE.References
@@
- Zhurko (2005) Zhurko, G. A. 2005, ivanovo, Russia
Appendix A Optimized geometries and calculated BE values
Geometry optimization of the ASW clusters and the BE values estimation of the 7 radicals are performed using DFT-B97X-D/6-311+G(d,p) level theory as stated in Section 2.
ASW | -H binding | Optimized energy⋆ | Relative energy | Optimized cluster |
clusters | sites avaiable | (Hartrees) | [Eref.= ASW(1)] (Kelvin) | structures |
ASW(1) | 7 | -1528.510023 | 0 | |
ASW(2) | 7 | -1528.499694 | 3262 | |
ASW(3) | 9 | -1528.505233 | 1513 | |
ASW(4) | 10 | -1528.494038 | 5048 | |
ASW(5) | 8 | -1528.499406 | 3353 | |
ASW(6) | 9 | -1528.494265 | 4976 | |
ASW(7) | 8 | -1528.508844 | 372 | |
ASW(8) | 9 | -1528.503436 | 2080 | |
ASW(9) | 8 | -1528.500329 | 3061 |
⋆ Optimized energy values correspond to the ZPE-corrected electronic energy in Hatree.
The sign indicates the stability of the structure; the larger it is, the more stable the optimized geometry.
-H binding sites | BE values (Kelvin) | Highest BE value (Kelvin) | ||
BSSE | BSSE | BSSE | BSSE | |
ASW(1) | ||||
1 | 2465 | 2162 | ||
2 | 2955 | 2674 | ||
3 | 1880 | 1500 | 3990 | 3643 |
4 | 3125 | 2816 | ||
5 | 3990 | 3643 | (S.F. ) | |
6 | 3116 | 2793 | ||
7 | 1662 | 1452 | ||
ASW(2) | ||||
1 | 1709 | 1495 | ||
2 | 1498 | 1265 | ||
3 | 2887 | 2553 | ||
4 | 4734 | 4296 | 4734 | 4296 |
5 | 2957 | 2630 | ||
6 | 3456 | 3130 | ||
7 | 4349 | 3914 | ||
ASW(3) | ||||
1 | 3410 | 3063 | ||
2 | 3435 | 2931 | ||
3 | 2939 | 2640 | ||
4 | 1683 | 1456 | ||
5 | 5580 | 5028 | 5580 | 5028 |
6 | 3906 | 3482 | ||
7 | 1758 | 1532 | ||
8 | 3346 | 2999 | ||
9 | 4492 | 4080 | ||
ASW(4) | ||||
1 | 3676 | 3331 | ||
2 | 3948 | 3481 | ||
3 | 3219 | 2888 | ||
4 | 2964 | 2508 | ||
5 | 1799 | 1573 | 4719 | 4261 |
6 | 4114 | 3654 | ||
7 | 3170 | 2867 | ||
8 | 4062 | 3579 | ||
9 | 4719 | 4261 | ||
10 | 3012 | 2687 | ||
ASW(5) | ||||
1 | 3582 | 3246 | ||
2 | 3909 | 3572 | ||
3 | 3150 | 2801 | ||
4 | 3642 | 3294 | ||
5 | 3561 | 3237 | 5397 | 4902 |
6 | 1914 | 1683 | ||
7 | 4285 | 3882 | ||
8 | 5397 | 4902 | ||
ASW(6) | ||||
1 | 2785 | 2320 | ||
2 | 3497 | 3168 | ||
3 | 3629 | 3300 | ||
4 | 4178 | 3774 | ||
5 | 4770 | 4191 | 4802 | 4366 |
6 | 4569 | 4046 | ||
7 | 3755 | 3399 | ||
8 | 4802 | 4366 | ||
9 | 4184 | 3780 | ||
ASW(7) | ||||
1 | 3798 | 3337 | ||
2 | 3468 | 3134 | ||
3 | 3525 | 3190 | ||
4 | 3340 | 3005 | ||
5 | 2605 | 2104 | 3798 | 3337 |
6 | 3560 | 3229 | ||
7 | 3315 | 2967 | ||
8 | 2972 | 2533 | ||
ASW(8) | ||||
1 | 4436 | 3986 | ||
2 | 1633 | 1428 | ||
3 | 2990 | 2586 | ||
4 | 3609 | 3270 | ||
5 | 3237 | 2875 | 4436 | 3986 |
6 | 3008 | 2686 | ||
7 | 3414 | 3077 | ||
8 | 2923 | 2618 | ||
9 | 3642 | 3226 | ||
ASW(9) | ||||
1 | 2803 | 2492 | ||
2 | 3236 | 2931 | ||
3 | 4972 | 4503 | ||
4 | 3132 | 2807 | ||
5 | 5216 | 4761 | 5216 | 4761 |
6 | 2751 | 2329 | ||
7 | 3560 | 3225 | ||
8 | 1746 | 1518 | ||
Total 75 sites | — | — | Average |
-H binding | BE values (Kelvin) | Highest BE value (Kelvin) | Scaled by | |||
sites of ASW(1) | BSSE | BSSE | BSSE | BSSE | 1.188 for BSSE | 1.177 for BSSE |
CH radical | ||||||
1 | 1837 | 1670 | ||||
2 | 1688 | 1530 | ||||
3 | 1610 | 1454 | ||||
4 | 1601 | 1446 | 1892 | 1737 | 2248 | 2044 |
5 | 31315∗ | — | ||||
6 | 1892 | 1737 | ||||
7 | 14746∗ | — | ||||
NH radical | ||||||
1 | 1441 | 1282 | ||||
2 | 1409 | 1258 | ||||
3 | 1363 | 1223 | ||||
4 | 1331 | 1182 | 2083 | 1898 | 2475 | 2234 |
5 | 2083 | 1898 | ||||
6 | 1569 | 1421 | ||||
7 | 1425 | 1280 | ||||
SH radical | ||||||
1 | 1661 | 1363 | ||||
2 | 1870 | 1610 | ||||
3 | 1010 | 866 | ||||
4 | 1839 | 1572 | 2416 | 2109 | 2870 | 2482 |
5 | 2416 | 2109 | ||||
6 | 1100 | 942 | ||||
7 | 1276 | 1131 | ||||
CN radical | ||||||
1 | 1180 | 1068 | ||||
2 | 1192 | 1086 | ||||
3 | 1121 | 1016 | ||||
4 | 1028 | 925 | 1192 | 1086 | 1416 | 1278 |
5 | 6652∗ | 6317∗ | ||||
6 | 1152 | 1047 | ||||
7 | 1146 | 1044 | ||||
NS radical | ||||||
1 | 2117 | 1735 | ||||
2 | 2038 | 1677 | ||||
3 | 1808 | 1473 | ||||
4 | 2316 | 1879 | 2606 | 2225 | 3096 | 2619 |
5 | 2606 | 2225 | ||||
6 | 1865 | 1513 | ||||
7 | 2462 | 2039 | ||||
NO radical | ||||||
1 | 686 | 510 | ||||
2 | 540 | 353 | ||||
3 | 895 | 591 | ||||
4 | 931 | 598 | 931 | 598 | 1106 | 704 |
5 | 743 | 574 | ||||
6 | 658 | 490 | ||||
7 | 384 | 259 |
-H | B97X-D / 6-311+G(d,p) | ||
binding sites | BSSE | BSSE | |
1 | 2465 | 2162 | 1.1401 |
2 | 2955 | 2674 | 1.1051 |
3 | 1880 | 1500 | 1.1253 |
4 | 3125 | 2816 | 1.1097 |
5 | 3990 | 3643 | 1.0952 |
6 | 3116 | 2793 | 1.1156 |
7 | 1662 | 1452 | 1.1446 |

### v2

@@ 3.3 Comparison of the BEs with previous studies
It is important to note that the small fluctuations observed in the calculated BEs compared to literature values can be attributed to several factors, including the specific computational method employed, the adopted ice model, and the particular binding sites investigated in this study. Open-shell species pose significant computational challenges due to multi-reference characteristics and the need to account for static and dynamic correlation effects. Accurately capturing the true energy of these species requires a sensitive and careful approach, often involving multi-reference wave function theories (e.g., the coupled cluster method) and perturbation methods (e.g., the Møller-Plesset method). Although these techniques can provide highly accurate results, they come with a substantial computational cost. The combination of the B97X-D functional and the 6-311+G(d,p) basis set offers a practical balance between accuracy and efficiency. The B97X-D functional is crucial for handling the complex electronic structures inherent to radicals and is a versatile tool for studying a wide range of systems. The 6-311+G(d,p) basis set, with its valence triple- quality and added polarization functions, further enhances the accuracy by providing a flexible and detailed description of electron density, especially where electron distribution is uneven due to unpaired electrons. This combination is not necessarily the best or most accurate option available. Still, it is a well-considered choice that delivers a high degree of accuracy while keeping computational costs relatively low. Furthermore, we believe that our methodology, with careful structural optimization and analysis, is robust enough to be applied to other radical species and, more generally, to species with challenging electronic structures.
@@ 4.1 Chemisorption
Some of our binding sites exhibit chemisorption for CH (a chemical bond is formed with one oxygen atom of a water molecule of the cluster). Figure A.3 (available on Zenodo) depicts the formation of CH2OH and CHOH2 by chemisorption. There is a possibility that the chemisorbed CH2OH could produce CH3OH via the Eley-Rideal mechanism by capturing an H atom from the gas phase. Alternatively, a chemisorbed H atom could react with CH2OH to form methanol. Recent experiments by Grieco et al. (2024) suggest the formation of H2O at a high temperature (approximately 85 K) by chemisorbed H and O. Additionally, Grieco et al. (2023) reported the formation of H2 up to a temperature as high as 250 K, suggesting the feasibility of hydrogenation reactions at higher temperatures than previously thought. Contrary to previous understandings of methanol formation through the surface hydrogenation of CO, Molpeceres et al. (2021) and Tsuge et al. (2023) suggest that before CO freeze-out, atomic carbon deposited on ASW could form H2CO. Tsuge et al. (2023) pointed out that at temperatures below 20 K, some of the accreted carbon atoms are quickly converted to formaldehyde. The remaining physisorbed atoms are partially changed to chemisorbed and hydrogenated to CH4. Physisorbed carbon atoms can diffuse at temperatures between 20 K and 30 K, promoting insertion reactions with other carbon atoms or species. Future observations with the James Webb Space Telescope (JWST) could provide additional details about the significance of chemisorption in forming complex organic molecules in the ice phase.
4.2 Implication of the computed BEs for the astrochemical model
@@
Recently, using Atacama Large Millimeter/submillimeter Array (ALMA) observations, Coutens et al. (2019) detected multiple transitions of HONO in the low-mass protostellar system IRAS 16293–2422. This system consists of three protostars, two located toward source A and one toward source B (Maureira et al., 2020). The identification by Coutens et al. (2019) specifically focuses on the protostar associated with source B. They also attempted to explain the abundances of HONO-related species and their ratios using a chemical model. However, they were unable to account for all the observed features. In their study, Coutens et al. (2019) estimated abundance ratios of HONO/NH2OH and HONO/HNO to be greater than or equal to and , respectively. Using their chemical model, they successfully explain the abundance of HONO, N2O, and NO2. However, their model resulted in an overproduction of NH2OH and HNO, leading to lower HONO/NH2OH () and HONO/HNO (0.01) ratios compared to the observed values. Our study suggests a weak binding of NO with the water substrate. It would be interesting to see how the estimated BEs would affect the modeled abundances of these species.
@@
We utilized the Rokko code Furuya et al. (2015, 2017) for our chemical modeling, which encompasses three phases: gas, icy grain surface, and homogeneous bulk ice mantle. Our active ice layers were limited to the top four monolayers. All layers, including the active ice layers, were considered part of the bulk ice mantle. The gas- and ice-phase networks are based on the work of Garrod (2013). The active layers undergo photodissociation and the photo-desorption of species. A constant reactive desorption of all the species by a factor of 0.01 is assumed, resulting in a single product. MD studies have determined that approximately 10% of the formation energy is retained after the formation of H2S from HS+H (Bariosco et al., 2024, and references therein). However, based on experimentally determined reaction desorption probabilities for H2S, HDS, and D2S (Oba et al., 2019; Furuya et al., 2022), we used a reactive desorption factor of 0.03. We also used a standard cosmic-ray ionization rate of s-1.
@@ 4.2.1 Dark cloud
For the dark cloud model, we started with a hydrogen number density, cm-3, a visual extinction () of 10 mag, and a gas and ice temperature of 10 K. The initial elemental abundances of species were taken from Garrod et al. (2017). The upper panel of Fig. 5 shows the time evolution of the radicals. Here, “G” indicates that all the BE sets are from Garrod (2013); “G+S” indicates that the BEs for the radicals are taken from this work and the other BEs from Garrod (2013). In this work we used a default diffusion-to-BE ratio of 0.4. Since no estimated BEs were available for HONO, H2NO2, and H3NO2, we used our estimated values of 2918 K, 4205 K, and 3044 K, respectively, in both “G” and “G+S.” Changes in the BE would have a minor effect on the gas-phase abundances of these radicals. However, they would have a major effect on the ice-phase abundances of these radicals, especially for the ice-phase NO and CH. The NO ice abundance is reduced by a factor of almost at years. Meanwhile, the ice-phase abundance of CH increases by a factor of 100. The reduction in the abundance of NO is due to the weak binding of NO obtained in this work ( K) compared to the earlier estimate (1600 K; Wakelam et al., 2017). The increase in the surface abundance of CH radicals is due to the higher estimated BE ( K) compared to previous work (Garrod, 2013, 925 K). The shaded lines on the graph indicate the uncertainty of our BE values from Table 2. Since NO has a very weak binding with the water substrate, its low uncertainty (about 94 K) in the BE estimation could cause a significant change in its ice-phase abundance (see the black shaded area in the top right panel of Fig. 5).
The lower panel of Fig. 5 presents the time evolution of the NO-related species. It clearly shows that with the present BE estimation of NO, we can expect a favorable formation of gas-phase HONO in a dark cloud. On the contrary, the formation of gas-phase NH2OH is heavily affected, which could be useful in explaining the observed HONO/NH2OH ratio in star-forming regions (Coutens et al., 2019) discussed in Sect. 4.2.2. The shaded regions are the results when the uncertainty in the BE calculations is taken into consideration. The obtained abundances and abundance ratio with HONO are noted in Table 4. Due to the weak binding of NO obtained in this study, we find a higher abundance of the ice-phase NO2 through the reaction , which results in a favorable formation of HONO-gr by . The formation of NH2OH-gr occurs through successive hydrogenation reactions: . When our BE value of NO is used, due to the high abundance of atomic oxygen on the grain compared to the atomic hydrogen, NO-gr is mainly channelized to form , resulting in a decrease in the formation of . On the contrary, it facilitates the HONO-gr formation. The lower panel of Fig. 5 presents the observed significant uncertainties in the ice-phase abundances of HNO and NH2OH due to the uncertainty surrounding the BE of NO ( 94 K).
@@
s When the BEs calculated here are used (otherwise those from Garrod 2013 are used).
@@ 4.2.2 Hot corino
We further examined a low-mass star-forming region to assess the impact of our proposed BEs on abundance estimations. We considered three stages: the isothermal collapsing stage, the warm-up stage, and the post-warm-up stage. During the isothermal collapsing stage, the cloud was allowed to collapse from a minimum density of cm-3 to a maximum density of cm-3 in 1 million years, with a constant gas temperature of 10 K and an ice temperature of 10 K. The initial abundance for this model was taken from the final abundance of the dark cloud model discussed in Sect. 4.2.1. The visual extinction in the first stage was allowed to vary with the density following the relation provided by Garrod & Pauly (2011), resulting in a variation of . In the warm-up stage, the density and visual extinction parameters were consistently set to their maximum values for an additional 1 million years, and the temperature gradually rose to a peak of K. During the post-warm-up stage, the parameters listed above remained at their maximum values for an additional years.
@@
The upper panel of Fig. 6 shows the time evolution of the abundances of gas-phase radicals. With the new BE value, the gas-phase abundance of NO (dashed black curve) decreases significantly during the warm-up stage. In the lower panel of Fig. 6, we show the abundances of NO and its related species separately. The uncertainties in the BE calculations are also considered here as discussed in the context of Fig. 5. In Table 4, we compare the peak abundances (peak value is taken beyond the isothermal stage) and the peak abundance ratio of some NO-related species obtained from our simulations with that observed by Coutens et al. (2019). With the new BEs (represented by the dashed curves), we observe a significant reduction in the abundances of NO, N2O, HNO, and NH2OH. Conversely, there is a surge in the abundance of HONO.
@@
NH2OH is primarily produced in the cold ice phase through the successive sequential hydrogenation of NO-gr, as discussed in Sect. 4.2.1, and it can sublimate to the gas phase after around years, depending on the BE (6806 K; Garrod, 2013). Due to the weak binding of NO, its abundance in the ice phase decreases as it forms NO2, leading to a decrease in NH2OH. We estimate a BE of 2918 K for HONO, so it is expected to be released earlier than NH2OH in the gas phase. However, an additional radical-radical pathway () is responsible for ice-phase HONO production in warmer regions. The low BE of NO obtained here further facilitates this radical-radical ice-phase reaction.
@@
In our study, with our estimated BE for NO, we obtain a lower abundance of NH2OH and HNO (see Fig. 6 and Table 4). We obtain peak abundance ratios of HONO/NH2OH and HONO/HNO of 143 and 3781, respectively, consistent with the observation. Table 4 shows that the other ratios from the observations seem to be between that found using the previous model (Coutens et al., 2019) and this model. This may indicate that the diffusion activation energy of NO was overestimated in the previous one and underestimated in the present.
@@ 5 Conclusions
This study aimed to calculate the BE of certain interstellar radicals on cluster surfaces of ASW. We employed a fully quantum chemical approach to examine a large number of binding sites for all the considered radicals using a suitable ASW cluster model. The significant findings of our work are as follows:
@@
We estimated the BEs of seven diatomic radicals using a pure quantum chemical approach with high-level water clusters and methods, and we discuss the implications of the computed BEs for astrochemical models.
@@
We carefully and extensively employed the BSSE correction using the CPC method. This correction is crucial, especially with small and medium-sized basis sets, for obtaining accurate and reliable estimates of the BEs between radical species and the ASW cluster.
@@
Our results suggest that, most of the time, we have physisorption, indicating weak bonding between radicals and the water molecule. However, in addition to physisorption, we find chemisorption of CH with the cluster in two cases. In these cases, a chemical bond is formed between CH and an oxygen atom of a water molecule. Figure A.3 (available on Zenodo) illustrates the production of CH2OH through the chemisorption of CH on a water substrate. This warrants the potential formation of methanol through hydrogenation in warmer regions. Future observations with the JWST could address this issue. Moreover, our estimated value of the BE (physisorption) of CH is 2-3 times higher than the available estimate. Hence, we suggest using these values for astrochemical models.
@@
We obtain a significantly lower BE value of NO (704 K), which is in excellent agreement with a very recent study (average value of 712 K with the ASW model; Martínez-Bachs et al., 2024) compared to that used earlier (1600 K). Our obtained BEs of NO can help explain the observed HONO/NH2OH and HONO/HNO ratio obtained in the B component of the low-mass protostellar binary IRAS 16293–2422.
Data availability
The data underlying this article (in the appendix) are made available under a Creative Commons Attribution license on Zenodo: doi:10.5281/zenodo.13388691.
@@ Acknowledgements.
The authors thank the referee for the valuable insights and constructive comments that significantly improved the manuscript. M.S. acknowledges financial support through the European Research Council (consolidated grant COLLEXISM, grant agreement ID: 811363). N.I. acknowledges the support of FONDECYT Regular Grant 1241193 and Vicerectoría de Investigación y Postgrado (VRIP). A.D. acknowledges the support of MPE for sponsoring a scientific visit to MPE.References